# Humanizing Machines: Rethinking LLM Anthropomorphism Through a Multi-Level Framework of Design

### v1

@@ Abstract
Yunze Xiao, Lynnette Hui Xian Ngthanks: Yunze Xiao and Lynnette Hui Xian Ng contributed equally to this paper, Jiarui Liu, Mona Diab Carnegie Mellon University {yunzex,huixiann,jiaruil5,mdiab}@andrew.cmu.edu
@@ 1 Introduction
Anthropomorphism is a purposeful design strategy that unlocks richer, more intuitive collaboration between humans and Artificial Intelligent (AI) systems. Developers equip large language models (LLMs) with relatable personalities Wang et al. (2024c); tse Huang et al. (2024b), emotional expressiveness Huang et al. (2024a), and context-sensitive social reasoning Nighojkar et al. (2025). These human-like cues allow users to converse with a system in familiar terms. This design principle builds user trust by reducing the cognitive effort required to interact with the system. Multi-modal extensions amplify this effect and create seamless and engaging interactions: speech synthesis models convey nuanced emotion Zhou et al. (2022), embodied agents navigate physical space Xie et al. (2025), and vision-based models interpret social scenes Mathur et al. (2025). These human-like agents have shown tangible benefits to our society: providing realistic training environments for education Ma et al. (2024), improving adherence and empathy for virtual healthcare consultations Wen et al. (2024), strengthening therapeutic alliances in psychiatry Wang et al. (2024a), or making legal reasoning tools accessible to non-experts Huang et al. (2023). In each scenario, carefully calibrated anthropomorphic cues bridge complex AI capabilities and human goals, enabling technology that feels / appears to be supportive, transparent, and responsive.
@@
This multimodal reconceptualization of anthropomorphism expands the research agenda, enabling scholars to systematically investigate the scenarios in which anthropomorphic design cues are beneficial. This conceptualization is critical because LLMs increasingly mediate experiences across sensitive and subjective domains, in which the trust and engagement of interpreters profoundly shape outcomes. The nuanced understanding of anthropomorphism is based on the interaction between user and model and focuses on the effectiveness of system design. This approach reduces mis-characterizations of system effects through over-reliance of cautionary narratives, and the deployment of human-like systems without adequate design foresight. Our definition thus provides a coherent frame for auditing human-imitative AI across varied modalities, bringing a new perspective to harness anthropomorphism responsibly and effectively.
@@ 3.4 Cognitive Cues (reasoning level signal)
Cognitive cues operationalized through system behaviors that suggest internal system deliberation or state modeling, even when no such process exists. LLMs reflect the cognitive dimension by alluding to thinking behavior with their outputs that simulate reasoning, reflection, or uncertainty. Such outputs serve as cognitive signals that prompt the interpreter to assign intelligence or thoughtfulness to the system.
@@
The implications of cognitive anthropomorphism are highly context-dependent. In the contexts of education and mental health, common artifacts are conversational chatbots. In these artifacts, high-level cognitive cues such as empathy expression, reflective revision, and logical reasoning, can enhance the interpreted competence of the artifacts. When interpreters view the artifacts as more thoughtful and emotionally aware, there is increased engagement and trust Gillath et al. (2021). However, in the context of web search and other task-specific applications, low-level cues that promote clarity and efficiency will suffice. Designers should thus calibrate the degree of cognitive cues to the context of artifact use in order to increase usability and user satisfaction.
@@ 5 Recommendations
Anthropomorphic design should not be treated as a one-size-fits-all solution; rather, it must be calibrated to the specific context in which an artifact is deployed. Designers should visualize and monitor the evolution of interpreter-artifact dynamics across repeated interactions, potentially using metrics such as trust calibration, emotional attribution, and perceived autonomy. Anthropomorphic intensity might need to start high to facilitate initial engagement but should ideally be adaptive. Design artifacts with sensitivity to cultural variation, as signals that build trust in one culture may cause discomfort in another. The cultural context should be treated as a dynamic design variable. The NLP community should engage more deeply in cultural adaptation, drawing inspiration from work such as Shiomi et al.. Examples are shown in Table 2.
@@ 6 Conclusion
Furthermore, anthropomorphic systems can entrench normative or cultural biases if behavioral and linguistic cues are not localized or participatory in design. This marginalizes underrepresented cultures or reinforce dominant interaction norms. Although our framework advocates for cross-cultural and context-sensitive cue calibration, more empirical research is needed to verify the effectiveness of such strategies across global deployments.

### v2

@@ Abstract
Yunze Xiao, Lynnette Hui Xian Ngthanks: Yunze Xiao and Lynnette Hui Xian Ng contributed equally to this paper, Jiarui Liu, Mona Diab Carnegie Mellon University {yunzex,huixiann,jiaruil5,mdiab}@cs.cmu.edu
@@ 1 Introduction
Anthropomorphism is a purposeful design strategy that unlocks richer, more intuitive collaboration between humans and Artificial Intelligent (AI) systems. Developers equip large language models (LLMs) with relatable personalities Wang et al. (2024c); tse Huang et al. (2024b), emotional expressiveness Huang et al. (2024a), and context-sensitive social reasoning Nighojkar et al. (2025); Liu et al. (2025). These human-like cues allow users to converse with a system in familiar terms. This design principle builds user trust by reducing the cognitive effort required to interact with the system. Multi-modal extensions amplify this effect and create seamless and engaging interactions: speech synthesis models convey nuanced emotion Zhou et al. (2022), embodied agents navigate physical space Xie et al. (2025), and vision-based models interpret social scenes Mathur et al. (2025). These human-like agents have shown tangible benefits to our society: providing realistic training environments for education Ma et al. (2024), improving adherence and empathy for virtual healthcare consultations Wen et al. (2024), strengthening therapeutic alliances in psychiatry Wang et al. (2024a), or making legal reasoning tools accessible to non-experts Huang et al. (2023). In each scenario, carefully calibrated anthropomorphic cues bridge complex AI capabilities and human goals, enabling technology that feels / appears to be supportive, transparent, and responsive.
@@
This multimodal reconceptualization of anthropomorphism expands the research agenda, enabling scholars to systematically investigate the scenarios in which anthropomorphic design cues are beneficial. This conceptualization is critical because LLMs increasingly mediate experiences in sensitive and subjective domains Ng et al. (2025), in which the trust and engagement of interpreters profoundly shape the outcomes. The nuanced understanding of anthropomorphism is based on the interaction between user and model and focuses on the effectiveness of system design. This approach reduces mis-characterizations of system effects through over-reliance of cautionary narratives, and the deployment of human-like systems without adequate design foresight. Our definition thus provides a coherent frame for auditing human-imitative AI across varied modalities, bringing a new perspective to harness anthropomorphism responsibly and effectively.
@@ 3.4 Cognitive Cues (reasoning level signal)
Cognitive cues operationalized through system behaviors that suggest internal system deliberation or state modelingXu et al. (2025), even when no such process exists. LLMs reflect the cognitive dimension by alluding to thinking behavior with their outputs that simulate reasoning, reflection, or uncertainty. Such outputs serve as cognitive signals that prompt the interpreter to assign intelligence or thoughtfulness to the system.
@@
The implications of cognitive anthropomorphism are highly context-dependent. In the contexts of education and mental health, common artifacts are conversational chatbots. In these artifacts, high-level cognitive cues such as empathy expression, reflective revisionYang et al. (2025), and logical reasoning, can enhance the interpreted competence of the artifacts. When interpreters view the artifacts as more thoughtful and emotionally aware, there is increased engagement and trust Gillath et al. (2021). However, in the context of web search and other task-specific applications, low-level cues that promote clarity and efficiency will suffice. Designers should thus calibrate the degree of cognitive cues to the context of artifact use in order to increase usability and user satisfaction.
@@ 5 Recommendations
Anthropomorphic design should not be treated as a one-size-fits-all solution; rather, it must be calibrated to the specific context in which an artifact is deployed. Designers should visualize and monitor the evolution of interpreter-artifact dynamics across repeated interactions, potentially using metrics such as trust calibration, emotional attribution, and perceived autonomy. Anthropomorphic intensity might need to start high to facilitate initial engagement, but should ideally be adaptive. Design artifacts with sensitivity to cultural variation, as signals that build trust in one culture may cause discomfort in another. The cultural context should be treated as a dynamic design variable. The NLP community should engage more deeply in cultural adaptation, drawing inspiration from work such as Shiomi et al.. Examples are shown in Table 2.
@@ 6 Conclusion
Furthermore, anthropomorphic systems can entrench normative or cultural biases if behavioral and linguistic cues are not localized or participatory in design. This marginalizes underrepresented cultures or reinforces dominant interaction norms. Although our framework advocates for cross-cultural and context-sensitive cue calibration, more empirical research is needed to verify the effectiveness of such strategies in global deploymentsDai and Xiao (2025).